OpenStudy (anonymous):

if a pig pen has a length that is two more than 4 times the width and the perimeter is 158, what are the dimensions of the pig pen

OpenStudy (anonymous):

what shape is the pen? rectangular?

OpenStudy (anonymous):

square

OpenStudy (anonymous):

no it cant be a square. it is a rectangle. anyway. do you know the formula for the perimeter of a rectangle?

OpenStudy (anonymous):

that all the info it gave

OpenStudy (anonymous):

okay dont worry about that. tell me if you know the formula for the perimeter of a rectangle

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay it is 2 * ( width + length)

OpenStudy (anonymous):

so lets call the length L and the width W .

OpenStudy (anonymous):

so what can you tell me about the relationship between L and W from your problem?

OpenStudy (anonymous):

2*6=12+158=170

OpenStudy (anonymous):

the problem states 2 more then 4 times the width and the perimeter is 158

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the length is 2 more than 4 times the width. If the width is W, then what is L?

OpenStudy (anonymous):

L=4W+2

OpenStudy (anonymous):

P=2(W+L) 158=2(W+4W+2) 158=2W+8W+4 158-4=2W+8W (154=10W) 1/10 15.4=W so thats it, the with of the pigpen is 15.4 to find the length, just use the formula and substitute the width with 15.4 P=2(W+L) 158=2 (15.4+L) 158= 30.8 +2L 158-30.8=2L (127.2=2L)1/2 63.6 = L so the length is 63.6. ^_^